# CogGuide: Human-Like Guidance for Zero-Shot Omni-Modal Reasoning

### v1

@@
Figure 1. Relative Percentage Improvements over Baselines by the Proposed Method on IntentBench, WorldSense, and Daily-Omni
@@ INTRODUCTION
Figure 2. Intent sketch guided zero-shot omni-modal reasoning component: three-module pipeline and information-entropy analysis
@@ 3.1 Experimental Settings
We evaluate the proposed method on multiple multimodal reasoning benchmark datasets, including the human-intent understanding task IntentBench, the audio–video collaborative analysis dataset WorldSense, and the daily-life scenario dataset Daily-Omni. Accuracy (%) is used as the evaluation metric to compare the baseline performance of the original models (without any strategy prompts) against the reasoning performance after introducing different combinations of modules. The method consists of three pluggable modules—Intent Perceiver, Strategy Generator, and Strategy Selector—and the final output is fed into a reasoning engine to produce answers (see Table 1, “Model Overview”).

### v2

@@
Figure 1. Relative Percentage Improvements over Baselines by the Proposed Method on IntentBench, WorldSense, and Daily-Omni.
@@ INTRODUCTION
Figure 2. Intent sketch guided zero-shot omni-modal reasoning component: three-module pipeline and information-entropy analysis.
@@ 2.4 A Unified Information-Theoretic Framework of Uncertainty Reduction
Algorithm 1: Pseudocode for intent sketch reasoning component
$\text{{Input }}X=(V,A,Q),\ k;\ \text{{Output }}y,R$
$I\leftarrow\mathrm{IntentPerceiver}(X)$
$S\leftarrow\mathrm{StrategyGenerator}(I,Q,k)=\{s_{1},\ldots,s_{k}\}$
$\text{{function}}\ \mathrm{POSTERIOR\_AND\_ENTROPY}(X,s)$
$\quad P_{Y}\leftarrow\mathrm{ReasoningEngine}.\mathrm{posterior}(Y\mid X,s)$
$\quad H\leftarrow\mathrm{Entropy}(P_{Y})$
$\quad\text{{return}}\ (P_{Y},H)$
$\mathrm{best\_H}\leftarrow+\infty;\ \mathrm{best\_s}\leftarrow\emptyset;\ \mathrm{post\_best}\leftarrow\emptyset$
$\text{{for each}}~s~\text{{in}}~S~\text{{do}}$
$\quad P_{Y},H\leftarrow\mathrm{POSTERIOR\_AND\_ENTROPY}(X,s)$
$\quad\text{if }H<\mathrm{best\_H}:$
$\qquad\mathrm{best\_H}\leftarrow H;\ \mathrm{best\_s}\leftarrow s;\ \mathrm{post\_best}\leftarrow P_{Y}$
end for
$s^{\ast}\leftarrow\mathrm{best\_s}$
$(R,y)\leftarrow\mathrm{ReasoningEngine}.\mathrm{solve\_with\_strategy}(X,s^{\ast})$
$\text{{return }}y,R$
Figure 3. Strategy-Conditioned Uncertainty Contraction.
@@ 3.1 Experimental Settings
We evaluate the proposed method on multiple multimodal reasoning benchmark datasets, including the human-intent understanding task IntentBench, the audio–video collaborative analysis dataset WorldSense, and the daily-life scenario dataset Daily-Omni. Accuracy (%) is used as the evaluation metric to compare the baseline performance of the original models (without any strategy prompts) against the reasoning performance after introducing different combinations of modules. The method consists of three pluggable modules—Intent Perceiver, Strategy Generator, and Strategy Selector—and the final output is fed into a reasoning engine to produce answers (see Table 1).